### v1

@@ Abstract
[Abstract]Simulation-based digital twins must provide accurate, robust and reliable digital representations of their physical counterparts. Quantifying the uncertainty in their predictions plays, therefore, a key role in making better-informed decisions that impact the actual system. The update of the simulation model based on data must be then carefully implemented. When applied to complex standing structures such as bridges, discrepancies between the computational model and the real system appear as model bias, which hinders the trustworthiness of the digital twin and increases its uncertainty. Classical Bayesian updating approaches aiming to infer the model parameters often fail at compensating for such model bias, leading to overconfident and unreliable predictions. In this paper, two alternative model bias identification approaches are evaluated in the context of their applicability to digital twins of bridges. A modularized version of Kennedy and O’Hagan’s approach and another one based on Orthogonal Gaussian Processes are compared with the classical Bayesian inference framework in a set of representative benchmarks. Additionally, two novel extensions are proposed for such models: the inclusion of noise-aware kernels and the introduction of additional variables not present in the computational model through the bias term. The integration of such approaches in the digital twin corrects the predictions, quantifies their uncertainty, estimates noise from unknown physical sources of error and provides further insight into the system by including additional pre-existing information without modifying the computational model.
@@
articletype: Special Issue Paperfootnotetext: Abbreviations: QoI, quantity of interest; SHM, structural health monitoring; GP, Gaussian process; OGP, orthogonal Gaussian process; KOH, Kennedy and O’Hagan.
@@ 1 Introduction
The extended availability of computational means in recent years has prompted the digitalization of a plethora of processes across industries. Nowadays, it is not unusual for practitioners, managers and engineers to have large quantities of data at their disposal to enhance their predictions and guide their decisions. Digital twins are key for efficiently harnessing such an amount of information. This is not different in the field of structural engineering and bridge management, where the utilization of digital twins of bridges provides immediate access to the past, current and future state of the structure. Nevertheless, one of the main challenges for its widespread implementation is the lack of homogenization and standardization in their definition and implementation1. Assuring the fidelity, accuracy and reliability of any particular digital twin and its predictions is therefore crucial in the absence of a common framework within which to evaluate its quality. This holds especially true for their applications to structures such as bridges, whose safety and integrity must be preserved under any condition. The objective of this work is to reliably quantify the discrepancy between model and observations in the context of stochastic simulation-based digital twins of bridges to assure their safe and robust deployment for real applications.
Despite considerable efforts, a concrete universally accepted definition of what is a digital twin is yet to be found2. Probably the most extended definition of a digital twin is the one given by Grieves and Vickers, who consider it as “a set of virtual information constructs that fully describe a potential or actual physical manufactured product from the micro atomic level to the macro geometrical level"3. However, this definition is too general for the context of this paper, which will focus on the representation and response of digital twins of bridges. We will alternatively consider the definition from Lu et al., where they refer to "a digital replica of physical assets, processes, and systems. Digital twins integrate artificial intelligence, machine learning, and data analytics to create living digital simulation models that can learn and update from multiple sources as well as represent and predict the current and future conditions of physical counterparts."4
In contrast to geometry-based digital twins, with a focus on geometrical representation of the information such as Building Information Models (BIM), and purely data-driven digital twins, which focus on sensor observations and monitoring, simulation-based digital twins complement the established knowledge from the system, present in a simulation model, with sensor observations. This allows for the calibration of a representation of the real state of the system that can be used to generate predictions on Quantities of Interest (QoI), while also posing further challenges. In particular, these predictions often present discrepancies when compared with sensor observations, which reduces their reliability5. To generate useful models that reflect their real counterparts, it is not only imperative to appropriately tune their associated parameters but also to improve the model where necessary. The update of the simulation model is often an iterative process guided by the available data and the requirements imposed on the model. As new data becomes available, the model may need to be updated or modified. Leveraging the available data and accounting for the shortcomings of the model pose a real challenges for keeping the simulation models of digital twins true to the observations and the real state of the system.
@@
One of the most popular methods to tune simulation models is using Bayesian inference approaches to estimate the latent parameters. This approach introduces information obtained from observations to update the models built following known physical laws. Due to its probabilistic nature, uncertainty quantification is also naturally included in the evaluation. The introduction of this uncertainty quantification implies a stochastic component in the predictions and definition of the digital twin. Additionally, as new measurements are obtained, the estimated model can be easily updated to reflect the new reality of the real system. This adaptability is obtained at the expense of often requiring reliable prior information on the potential values of the model parameters and needing numerous evaluations of usually very costly models. In structural and civil engineering, model updating using Bayesian inference has been applied to numerous aspects of the design and monitoring of structures. Material laws and constitutive relations are modified to reflect the behaviour of the real system6, 7, 8, 9, 10. The models that represent the structure are updated to match the observations11, 12, 13. In cases where the structural model presents a heterogeneous material behaviour, Bayesian inference methods are used to obtain the parameters of the representation of the material variability 14. For the case of bridges, SHM systems are calibrated following this methodology 15, bridge model parameters are obtained using data from measurement campaigns 16 and digital twins are tuned to represent the response of a bridge in real time17.
@@
Three main challenges arise when implementing a Bayesian framework for digital twins of structures such as bridges: data availability, model complexity and compatibility in the implementation18. Typically, large quantities of data are required for an accurate model updating, but sensor systems, despite providing continuous data streams over time, are usually limited to sparsely distributed observation points. Therefore, the fidelity of the prediction depends to a large extent on the data richness and quality. On the other hand, complex models are often unavoidable in the case of simulation-based digital twins. Surrogate models are then introduced as substitutes in time-sensitive applications or to reduce the computational cost for the Bayesian inference at the expense of introducing further uncertainties5, 19, 20. Finally, a large selection of software tools has been developed in recent years that allow for simplified implementations of Bayesian frameworks in larger digital set-ups. Examples of them are emcee21, PyMC22, stan23, UQpy24 or probeye25. In this project, the model updating schemes are implemented using probeye, which is an open-source package for Python developed to this end. However, no matter how finely tuned or complex a model is, it will still be limited by its own definition based on a set of prescribed assumptions. The reality, in contrast, is infinitely complex and will never be fully described by a computational model26. In the particular case of models updated in a Bayesian framework, the dependency on the training dataset often leads to overfitted and overconfident models that do not reflect the real system 27. One of the main targets of this work is precisely the quantification of those model discrepancies rooted in deficiencies in the model definition.
For a proper evaluation of the fidelity and reliability of a model, we find it essential to quantify explicitly the discrepancy between the model and the real system. For the purpose of this paper, we will define as model inadequacies all deviations in the behaviour of any subcomponent of the computational model when compared with a hypothetical real system. This applies to deviations in the value of the model parameters, boundary and initial conditions, variations in the constitutive law or directly the presence of effects in the real system that are not captured by the model. We will define as model bias or model discrepancy indistinctively the additive correction term necessary to compensate for the difference in outputs of the computational model and real system. This is coherent with the definition of model bias as typically implemented following Kennedy and O’Hagan’s framework28. In this paper, we will address the quantification of such a model bias paying special attention to its interpretation, leaving the model inadequacies at subcomponent level out of scope.
@@
The inclusion of model bias in the inverse problem formulation not only allows quantifying the discrepancy between model and observations, but is also able to improve the inference procedure29. However, the introduction of an overly flexible term to fully capture the model bias incurs potential identifiability issues 30. Throughout the years, several approaches have tried to tackle the identifiability issue, some of which are promising alternatives for their implementation in a digital twin. Brynjardóttir and O’Hagan31 highlight the impact of the choice of priors for the bias distribution, outlining the importance of elicitation and expert knowledge in their definition. Modularized versions of Kennedy and O’Hagan’s framework32 attempt to separate the effects of calibrating the model parameter from the bias. Successful examples are its application to energy systems33 and building energy transference estimation34. Alternatively, embedded approaches associate additively the bias term with the corresponding latent parameter35 or subcomponents of the system36, and have successfully been applied to composite rotor blades37, rocket ignition mechanisms35, jet turbulences 38 and healthcare policy models36, 39. Finally, modelling the bias term as Orthogonal Gaussian Processes (OGPs) 40 has been proposed as an alternative solution to prevent the identifiability problem between latent and bias parameters while recovering control over the optimality of the process. In this paper, we will compare the suitability of a modularized implementation of the model bias identification framework and the inclusion of OGPs. These two methods are the least intrusive for identifying the model bias, especially in comparison to embedded approaches, this makes them very attractive for digital twins, where modifications of the model are not always possible. To our knowledge, none of these approaches has been implemented or evaluated in the context of digital twins of bridges. Correlated or non-additive bias structures are out of the scope of this paper due to the increased complexity and amount of system information required for their implementation, which may not necessarily be available.
@@
As previously mentioned, the aim of this study is to implement, assess and extend a model bias identification methodology to be applied to digital twins of bridges, mitigating some of the problems present in Bayesian updating. In particular, we will show that including a bias term corrects the predictions, enhances the model and provides more reliable uncertainty thresholds. Three different approaches will be evaluated: the classical Bayesian inference approach without bias, a modularized version of Kennedy and O’Hagan’s framework32, and a modification of the same framework to include OGPs in the model bias definition40. To further take advantage of the particularities of Digital Twins, two novel extensions based on the bias formulation are developed in this work: the quantification of noise from physical sources in the systems response and the extension of the input domain with additional variables. On the one hand, the sensor’s network of the Digital Twin often provides a continuous stream of data that is susceptible to present variations due to unaccounted sources. Due to the existence of discrepancies between model and reality, the variance produced by such sources cannot be quantified adequately in many cases, as the model bias is often included in the noise estimation, leading to inaccurate results. The introduction of noise-aware kernels in the bias term enables such a quantification. On the other hand, Digital Twins present at their disposal a plethora of information sources that can be used to enrich the model. However, adapting the computational model to include them typically implies increasing its complexity and it is often not possible. Our proposed extension of the model bias term includes the desired additional variables non-intrusively, providing insight into their influence on the model. Both extensions will be implemented for the aforementioned approaches of interest in the context of Digital Twins of bridges.
@@
The different approaches will be presented in Section 2 and applied in Section 3. First, in Section 3.1, we will use a simple 1D example to compare how accurate the inferred parameters with each of the approaches are and evaluate their capacity to quantify the model bias. Then, in Section 3.2, we will analyze in a stochastic cantilever beam system based on a cantilever bridge their ability to quantify the noise from non-prescribed sources and the possibility to include heteroscedastic noise structures with physical meaning. Finally, in Section 3.3, we will implement each of the approaches to a simplified illustrative model based on a real bridge. We will prove how it is possible to include additional information external to the model, such as a temperature variation, to compensate for the bias in the model related to it. The comparisons will always be done under the optic of their potential application to a digital twin. Section 4 closes with the conclusions and potential extensions.
@@ 2 Methodology
The computational model $f$ is defined as a $q$-real valued function over a $p$-dimensional domain such that $f:\mathbb{R}^{p}\to\mathbb{R}^{q}$. Ultimately, the goal is to obtain a computational model $f$ that behaves as the real system $z$, represented as an analogously defined function. The computational model $f$ takes as input a set of system coordinates $x\in\mathbb{R}^{p}$, is defined by a set of latent parameters $\theta$ and a set of prescribed parameters $\lambda$ and produces an output $u\in\mathbb{R}^{q}$. In this paper, $x$ denotes the input variable of $f$, $\mathbf{x}$ denotes a discrete realization of $x$ represented as a $p$-dimensional vector, $\hat{\mathbf{x}}$ denotes a sample of $n$ discrete realizations $\mathbf{x}$ represented as a $n\times p$ matrix, and $\mathbf{x}_{i}$ denotes the $i$-th entry of $\hat{\mathbf{x}}$. In practice, $\mathbf{x}$ corresponds to the input coordinates at which one observation takes place, e.g. a sensor’s spatial coordinates, and $\hat{\mathbf{x}}$ the set of input coordinates where each of the observations takes place. Following the same notation, $\hat{\mathbf{y}}$ represents the set of $n$ observations employed to fit the parameters of $f$. These observations are assumed to be produced by a generative process $y$ defined as a $q$-real valued function. The observations may not only depend on $x$ but also admit the coordinates $\eta\in\mathbb{R}^{d}$ as an extended input not included in $f$, with its associated set of $n$ realizations $\hat{\bm{\eta}}$. An example of such variable $\eta$ may be a set of time or temperature coordinates that are known for each observation but not considered by the computational model $f$. Then, the generative process is defined as $y(x,\eta):\mathbb{R}^{p+d}\to\mathbb{R}^{q}$. These observations are generally provided by physical sensors with a prescribed measurement noise $\epsilon_{\text{meas}}$, which corresponds to the difference between observations and real system $z$. Analogously, the real system is defined as a function $z(x,\eta):\mathbb{R}^{p+d}\to\mathbb{R}^{q}$. In a general continuous way, we can formulate the problem as
$y(x,\eta)=z(x,\eta)+\epsilon_{\text{meas}}(x,\eta)=f(x,\theta,\lambda)+\epsilon_{\text{model}}(x,\eta)+\epsilon_{\text{meas}}(x,\eta)$ (1)
where $\epsilon_{\text{model}}$ is an additional error term indicating the discrepancy between the computational model and reality, necessary to achieve the equality. For example, such a formulation can be applied to the calibration of the Young’s modulus of a linear elastic cantilever beam under self-weight. We consider $x$ as the set of spatial coordinates that identify the different points of the beam. The cantilever beam deformation is governed by its Young’s modulus $E$, which would be the latent parameter $\theta$ and other material and geometry parameters such as length, density, Poisson’s ratio, etc., that are encoded in $\lambda$. The set of equations of linear elasticity that for a given $E$ provides the deformation of this specific beam at a position $x$ is the computational model $f$. The real deformations of the cantilever beam are defined by $z$, which follows an infinitely complex law. Moreover, $z$ does not only depend on $x$, but on other parameters $\eta$, such as the temperature of the beam. The discrepancy between $z$ and $f$ due to the necessary simplifications in the system are $\epsilon_{\text{model}}$, which can explicitly include the temperature dependency $\eta$, present in $z$ but not in $f$. Finally, $z$ would be observed through some type of sensor, image correlation for example, that introduces a noise $\epsilon_{\text{meas}}$ in the measurement $y$, which may be different for every point and at every temperature.
During this paper, the parameters $\lambda$ will not be explicitly included in the formulation to avoid excessive notation. If we consider only a discrete set of $n$ observations, then the relationship is transformed to Equation 2:
$\hat{\mathbf{y}}=f(\hat{\mathbf{x}},\theta)+\epsilon_{\text{model}}(\hat{\mathbf{x}},\hat{\bm{\eta}})+\epsilon_{\text{meas}}(\hat{\mathbf{x}},\hat{\bm{\eta}}).$ (2)
@@
An important modelling choice resides in the structure of the error terms $\epsilon_{\text{model}}$ and $\epsilon_{\text{meas}}$. The term $\epsilon_{\text{meas}}$ defines exclusively the prescribed noise from the sensor. It is generally provided by the manufacturer in the cases of physical sensors or prescribed in the case of virtual ones. Noise will be added to the covariance matrix of the predictions from $z(x,\eta)$. For simplicity, it will be considered uncorrelated Gaussian noise with a prescribed amplitude $\sigma$. The term $\epsilon_{\text{model}}$ corresponds to the manifestation of the model discrepancy to the observed quantities, which is the object of this study. The reduction of $\epsilon_{\text{model}}$ is a model identification problem, usually following an iterative approach. In this context, changes to the model are introduced to better match the available information and minimize the discrepancy with the real system. This model update must then consider the available data to identify the fitting parameters and guide potential further improvements in subsequent iterations.
@@
The objective of the parameter identification approaches for one of such iterations is to find parameters $\theta$ from Equation 2 such that the deviation of the computational from the physical system is minimized. This deviation can be quantified in many different ways depending on what is understood as the “best-fitting" parameters $\theta$. A classical metric for such an approach is the Mean Squared Error (MSE), under which the problem reduces to solving:
$\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}\left(\mathbf{y}_{i}-f(\mathbf{x}_{i},\theta)\right)^{2}.$ (3)
@@
This concept can be extended to a functional analysis framework, allowing to define losses over the functions $y$ and $f$ instead of its observations. The main loss considered in this paper is the $L^{2}$ loss, although other options such as $W_{1}^{2}$ can be implemented as well. The $L^{2}$ loss is generated from the homonymous norm in its associated Hilbert space, being analogous to the MSE in a functional sense40:
$L_{L^{2}}\left\{y(\cdot)-f(\cdot,\theta)\right\}:=\int_{X}(y(x)-f(xi,\theta))^{2}\text{d}x=\left\|y(\cdot)-f(\cdot,\theta)\right\|^{2}_{L^{2}}$ (4)
@@
where $X$ is the domain of evaluation of the norm and no additional variables $\eta$ are considered. The evaluation of such a loss in its continuous form is not possible in practice, hence discretization is required. The main difference with MSE is that $L^{2}$ depends on the discretization of the loss itself while MSE relies exclusively on the observations, which provides more flexibility. This is particularly useful for the case that sparse observations do not correspond with the points of interest, which can make use of alternative discretizations to cover them.
@@
The inverse problem minimizes any of the previous losses varying $\theta$. A priori, these metrics would require extensive knowledge about both the generative and the computational model, but they would achieve the best possible fit under a given set of conditions. If the samples represent sufficiently well the domain of the generative process and the predictions are generated with enough information from the computational model, MSE and $L^{2}$ losses tend to the same value.

### v2

@@ Abstract
[Abstract]Simulation-based digital twins must provide accurate, robust and reliable digital representations of their physical counterparts. Quantifying the uncertainty in their predictions plays, therefore, a key role in making better-informed decisions that impact the actual system. The update of the simulation model based on data must be then carefully implemented. When applied to complex structures such as bridges, discrepancies between the computational model and the real system appear as model bias, which hinders the trustworthiness of the digital twin and increases its uncertainty. Classical Bayesian updating approaches aiming to infer the model parameters often fail at compensating for such model bias, leading to overconfident and unreliable predictions. In this paper, two alternative model bias identification approaches are evaluated in the context of their applicability to digital twins of bridges. A modularized version of Kennedy and O’Hagan’s approach and another one based on Orthogonal Gaussian Processes are compared with the classical Bayesian inference framework in a set of representative benchmarks. Additionally, two novel extensions are proposed for such models: the inclusion of noise-aware kernels and the introduction of additional variables not present in the computational model through the bias term. The integration of such approaches in the digital twin corrects the predictions, quantifies their uncertainty, estimates noise from unknown physical sources of error and provides further insight into the system by including additional pre-existing information without modifying the computational model.
@@
articletype: Special Issue Paperfootnotetext: Abbreviations: SHM, structural health monitoring; GP, Gaussian process; OGP, orthogonal Gaussian process; KOH, Kennedy and O’Hagan.
@@ 1 Introduction
The extended availability of computational means in recent years has prompted the digitalization of a plethora of processes across industries. Nowadays, it is not unusual for practitioners, managers and engineers to have large quantities of data at their disposal to enhance their predictions and guide their decisions. Digital twins are key for efficiently harnessing such an amount of information. This is not different in the field of structural engineering and bridge management, where the utilization of digital twins of bridges provides immediate access to the past, current and future state of the structure. Despite recent attempts on its classification1, one of the main challenges for its widespread implementation is the lack of homogenization and standardization in their definition and implementation2. Assuring the fidelity, accuracy and reliability of any particular digital twin and its predictions is therefore crucial in the absence of a common framework within which to evaluate its quality. This holds especially true for applications to structures such as bridges, whose safety and integrity must be preserved under any condition. The objective of this work is to reliably quantify the discrepancy between model and observations in the context of stochastic simulation-based digital twins of bridges to assure their safe and robust deployment for real applications. The main focus will therefore be on those digital twins that use models based on physics and incorporate real sensor data from a structure, e.g. a bridge, representing its current state. They are the so-called "dynamic system models" 3, 4, 5, and their goal is to analyze and monitor the actual state of the structure, for which their predictive power is crucial. Nevertheless, the techniques here implemented are designed to be applicable to any system where a model must replicate a set of observations. Therefore, the additional challenges tied to the dynamic nature of the predictions, in a digital twin 5, e.g. state-transition uncertainty quantification or real-time computation, remains out of scope.
In contrast to geometry-based digital twins, with a focus on geometrical representation of the information such as Building Information Models (BIM), and purely data-driven digital twins, which focus on sensor observations and monitoring, simulation-based digital twins complement the established knowledge from the system, present in a simulation model, with sensor observations. Following the usual formalization6, 7, the simulation or computational model can be represented by a function $f:\mathbb{R}^{p}\to\mathbb{R}^{q}$ that depends on a set of parameters $(\bm{\theta},\bm{\lambda})$ that control its behaviour and takes as input a vector of coordinates or variables $\mathbf{x}\in\mathbb{R}^{p}$. We differentiate the vector of latent parameters $\bm{\theta}\in\mathbb{R}^{s}$, whose values are unknown beforehand and must be obtained using the observations, from the vector of prescribed parameters $\bm{\lambda}\in\mathbb{R}^{t}$, whose values are set by the modeller beforehand. The real system $z$ is represented by an analogous function that does not depend on $(\bm{\theta},\bm{\lambda})$, as it reflects an infinitely-complex reality without a known parametrization. If we consider a vector of sensor observations $\mathbf{y}_{i}$ for $\mathbf{x}_{i}$ from the set of all $n$ collected observations $\hat{\mathbf{y}}$ taken at the set of coordinate vectors $\hat{\mathbf{y}}$, the realization is obtained from the statistical model as
$\mathbf{y}_{i}=z(\mathbf{x}_{i})+\epsilon_{\text{meas}}(\mathbf{x}_{i})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall i\in[1,...,n]$ (1)
where $\epsilon$ represents a measurement error or noise that may depend on $\mathbf{x}_{i}$. It is generally not possible to access $z$, therefore $f$ is used in its place and the main objective is to obtain the parameters $\bm{\theta}$ best represent the real system. In particular, these predictions often present discrepancies when compared with sensor observations, which reduces their reliability8. Considering the computational model $f$, the statistical model results in
$\mathbf{y}_{i}=f(\bm{\theta},\mathbf{x}_{i})+\epsilon_{\text{model}}(\mathbf{x}_{i})+\epsilon_{\text{meas}}(\mathbf{x}_{i})\leavevmode\nobreak\ \leavevmode\nobreak\ \forall i\in[1,...,n]$ (2)
where $\epsilon_{\text{model}}$ refers the so-called model discrepancy or model bias. The function $y(x)$ will refer to the generator model analogous to $z$ that produces the discrete set of observations $\hat{\mathbf{y}}=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{n}\}$.
@@
One of the most popular methods to obtain $\theta$ is using Bayesian approaches. Due to its probabilistic nature, uncertainty quantification is also naturally included in the evaluation. The introduction of this uncertainty quantification implies a stochastic component in the predictions and definition of the digital twin. Additionally, as new measurements are obtained, the estimated model can be easily updated to reflect the new reality of the real system. This adaptability is obtained at the expense of often requiring reliable prior information on the potential values of the model parameters and needing numerous evaluations of usually very costly models. For the case of bridges, Structural Health Monitoring (SHM) systems are calibrated following this methodology 9, bridge model parameters are obtained using data from measurement campaigns 10 and digital twins are tuned to represent the response of a bridge in real time11.
@@
Three main challenges arise when implementing a Bayesian framework for digital twins of structures such as bridges in a industrial application: data availability, model complexity and the compatibility between different software pieces and the interfaces that allow the exchange of information between the modules12. Typically, large quantities of data are required for an accurate model updating, but sensor systems of bridges, despite providing continuous data streams over time, are usually limited to sparsely distributed observation points due to installation and maintenance costs of sensors, as in 13 for example. Despite attempts to increase the available information through data assimilation approaches 14, 15, 16, the fidelity of the prediction still depends to a large extent on the data richness and quality. On the other hand, complex models are often unavoidable in the case of simulation-based digital twins. Surrogate models are then introduced as substitutes in time-sensitive applications or to reduce the computational cost for the Bayesian calibration at the expense of introducing further uncertainties17, 18, 19. The quantification of both surrogate and model discrepancy results critical for such cases. Finally, a large selection of software tools has been developed in recent years that allow for simplified implementations of Bayesian frameworks in larger digital set-ups. Examples of them are emcee20, PyMC21, stan22, UQpy23 or probeye24. In this project, the model updating schemes are implemented using probeye, which is an open-source package for Python developed to this end. However, no matter how finely tuned or complex a model is, it will still be limited by its own definition based on a set of prescribed assumptions. The reality, in contrast, is infinitely complex and will never be fully described by a computational model25. In the particular case of models updated in a Bayesian framework, the dependency on the training dataset often leads to overfitted and overconfident models that do not reflect the real system 26. One of the main targets of this work is precisely the quantification of those model discrepancies rooted in deficiencies in the model definition.
For a proper evaluation of the fidelity and reliability of a model, we find it essential to quantify explicitly the discrepancy between the model and the real system. For the purpose of this paper, we tackle the challenges presented by the model discrepancy $\epsilon_{\text{model}}$ as described in Equation 2 implemented following Kennedy and O’Hagan’s framework27 as an additive term to the model output, leaving the discrepancy at subcomponent level out of scope.
@@
The inclusion of model bias in the inverse problem formulation not only allows quantifying the discrepancy between model and observations, but is also able to improve the inference procedure28. However, the introduction of an overly flexible term to fully capture the model bias incurs potential identifiability issues 29. Throughout the years, several approaches have tried to tackle the identifiability issue, some of which are promising alternatives for their implementation in a digital twin. Brynjardóttir and O’Hagan30 highlight the impact of the choice of priors for the bias distribution, outlining the importance of elicitation and expert knowledge in their definition. Modularized versions of Kennedy and O’Hagan’s framework31 attempt to separate the effects of calibrating the model parameter from the bias. Successful examples are its application to energy systems32 and building energy transference estimation33. Non-additive bias structures e.g. using a multiplicative model, are investigated in 34, 35, 36 combined with the inclusion of discrepancy terms under Kennedy and O’Hagan’s framework in a way that minimizes identifiability problems, which has led to promising results in the field of hydraulics and nuclear engineering. Alternatively, embedded approaches associate additively the bias term with the corresponding latent parameter37 or subcomponents of the system38, and have successfully been applied to composite rotor blades39, rocket ignition mechanisms37, jet turbulences 40 and healthcare policy models38, 41. Finally, modelling the bias term as Orthogonal Gaussian Processes (OGPs) 42 has been proposed as an alternative solution to prevent the identifiability problem between latent and bias parameters while recovering control over the optimality of the process. In this paper, we will compare the suitability of a modularized implementation of the model bias identification framework and the inclusion of OGPs. These two methods are the least intrusive for identifying the model bias, especially in comparison to embedded approaches. This makes them very attractive for digital twins, where modifications of the model are not always possible. To our knowledge, none of these approaches has been implemented or evaluated in the context of digital twins of bridges. Correlated or non-additive bias structures are out of the scope of this paper due to the increased complexity and amount of system information required for their implementation, which may not necessarily be available.
@@
The aim of this study is to implement, assess and extend a model bias identification methodology to be applied to digital twins of bridges, mitigating some of the problems present in Bayesian updating. In particular, we will show that including a bias term corrects the predictions, enhances the model and provides more reliable uncertainty thresholds. Three different approaches will be evaluated: the classical Bayesian inference approach without bias, a modularized version of Kennedy and O’Hagan’s framework31, and a modification of the same framework to include OGPs in the model bias definition42. To further take advantage of the particularities of Digital Twins, two extensions are developed in this work: the quantification of noise from physical sources in the systems response and the extension of the input domain with additional variables. On the one hand, the sensor’s network of the Digital Twin often provides a continuous stream of data that is susceptible to present variations due to unaccounted sources. For example, it is not unusual that the deformation sensors installed in bridge monitoring systems have to be compensated due to temperature changes. If the computational model does not take into consideration the temperature, discrepancies between simulations and reality appear 43. It is not always possible to identify the source of such discrepancies, which very often appear in the form of noise that must be quantified. The introduction of noise-aware kernels in the bias term enables such a quantification in those cases that lack information of the noise’s origin. However, Digital Twins present at their disposal a plethora of documented information sources that can be used to enrich the model. Nevertheless, adapting the computational model to include them typically implies increasing its complexity. Our proposed extension of the model bias term includes the desired additional variables non-intrusively, providing insight into their influence on the model. It is possible, for example, to use the temperature data available for the bridge through its inclusion in the model bias term without needing to modify the computational model. Both extensions will be implemented for the aforementioned approaches of interest in the context of Digital Twins of bridges.
@@
The different approaches will be presented in Section 2 and applied in Section 3. In Section 3.1, we will use a simple 1D example to compare how accurate the inferred parameters with each of the approaches are and evaluate their capacity to quantify the model bias. Then, in Section 3.2, we will analyze in a stochastic cantilever bridge their ability to quantify the noise from non-prescribed sources and the possibility to include heteroscedastic noise structures with physical meaning. Finally, in Section 3.3, we will implement each of the approaches to a simplified illustrative model based on a real bridge. We will prove how it is possible to include additional information external to the model, such as a temperature variation, to compensate for the bias in the model. Section 4 closes with the conclusions and potential extensions.
@@ 2 Methodology
The methodology takes as base the formulation from Equation 2. An important modelling choice resides in the structure of the error terms $\epsilon_{\text{model}}$ and $\epsilon_{\text{meas}}$. The term $\epsilon_{\text{meas}}$ defines exclusively the prescribed noise from the sensor. It is generally provided by the manufacturer in the cases of physical sensors or prescribed in the case of virtual ones. Noise will be added to the covariance matrix of the predictions from $z(x,\eta)$. For simplicity, it will be considered uncorrelated Gaussian noise with a prescribed amplitude $\sigma$. The term $\epsilon_{\text{model}}$ corresponds to the manifestation of the model discrepancy to the observed quantities, which is the object of this study. The reduction of $\epsilon_{\text{model}}$ is a model identification problem, usually following an iterative approach. In this context, changes to the model are introduced to better match the available information and minimize the discrepancy with the real system. This model update must then consider the available data to identify the fitting parameters and guide potential further improvements in subsequent iterations.
Given a distance $L\left\{y(\cdot)-f(\cdot,\theta)\right\}$ between generating and computational models, the objective of the calibration approaches in a frequentist sense is to find the value of the estimated optimum $\bm{\theta}^{*}$ that is solution for the problem
$\bm{\theta}^{*}=\underset{\bm{\theta}}{\arg}\min L\left\{y(\cdot)-f(\cdot,\bm{\theta})\right\}.$ (3)
@@
This deviation can be quantified in many different ways depending on what is understood as the “best-fitting" parameters $\theta$. A classical metric for such an approach is the Mean Squared Error (MSE), under which the problem reduces to solving:
$\min_{\theta}\frac{1}{n}\sum_{i=1}^{n}\left(\mathbf{y}_{i}-f(\mathbf{x}_{i},\bm{\theta})\right)^{2}.$ (4)
@@
This concept can be extended to a functional analysis framework, allowing to define losses over the functions $y$ and $f$ instead of its observations. The main loss considered in this paper is the $L^{2}$ loss, although other options such as $W_{1}^{2}$ can be implemented as well. The $L^{2}$ loss is generated from the homonymous norm in its associated Hilbert space, being analogous to the MSE in a functional sense42:
$L_{L^{2}}\left\{y(\cdot)-f(\cdot,\bm{\theta})\right\}:=\int_{X}(y(x)-f(x,\bm{\theta}))^{2}\text{d}x=\left\|y(\cdot)-f(\cdot,\bm{\theta})\right\|^{2}_{L^{2}}$ (5)
@@
where $X$ is the domain of evaluation of the norm. The evaluation of such a loss in its continuous form is not possible in practice, hence discretization is required. The main difference with MSE is that $L^{2}$ depends on the discretization of the loss itself while MSE relies exclusively on the observations, which provides more flexibility. This is particularly useful for the case that sparse observations do not correspond with the points of interest, which can make use of alternative discretizations to cover them.
@@
A priori, these metrics would require extensive knowledge about both the generative and the computational model, but they would achieve the best possible fit under a given set of conditions. If the samples represent sufficiently well the domain of the generative process and the predictions are generated with enough information from the computational model, MSE and $L^{2}$ losses tend to the same value. These frequentist approaches do not inherently provide an uncertainty quantification nor leverage prior information on the parameters. These are desirable qualities in digital twins, therefore Bayesian approaches are favoured.